OpenStudy (anonymous):

Two planes left an airport at noon. One flew east and another flew west at twice the speed. After 3 hours the planes were 2700 mi apart. How fast was each plane flying?

OpenStudy (anonymous):

Something tells me it involves a formula like this: x+2y=2700 Or is it x+2x=2700? Thanks.

OpenStudy (anonymous):

I have no clue.....

OpenStudy (anonymous):

speed = Distance/ rate

OpenStudy (anonymous):

So S = 2700/r?

OpenStudy (anonymous):

900 mph 2700/3

OpenStudy (anonymous):

go it?

OpenStudy (anonymous):

got*

OpenStudy (anonymous):

Says in my answer sheet that it's 300 and 600 mph

OpenStudy (anonymous):

waitttttt hollup i overlooked twice as fast lemme redo it playa

OpenStudy (anonymous):

@aliboo997 @jim_thompson5910

OpenStudy (anonymous):

Hmm....searching google, has some good info

OpenStudy (anonymous):

Ok got it...was very simple. Don't know why I didn't get it sooner. Thanks for your help

OpenStudy (anonymous):

2x+x=900 3x=900 x=300 Answer: 300 and 600 For those who need help

OpenStudy (anonymous):

why is it 2x+x?

OpenStudy (anonymous):

Because if you solve for x, you can easily solve for 2x. So it would be best to combine the speeds of the two planes.